An organic compound contains 8% of sulphur . The minimum molar mass of the substance is 
a) 500 g/mol
b) 400 g/mol
c) 800 g/mol 
d) 600 g/mol

Grade:12th pass

2 Answers

Arun
4 years ago

Atomic mass of Sulphur is 32 AMU

if 8% of the molecular mass is sulphur or 32 AMU

total molecular mass = 32x100/8 AMU = 400 AMU

Khimraj
4 years ago

8 g S present in 100 g substance, So 32 g of S present in…….? =32*100/8=400